The ratio of the ages of a mother and daughter is 5 : 1
The mother is 30 years old at present.
In how many years will the age ratio be 3 : 1?

Answer by Maths68(1474)

Let
Present age of daughters = d
Present age of mothers age = m


The ratio of the ages of a mother and daughter is 5 : 1
m:d=5:1
m/d=5
m=5d
We have been told the mother is 30 years old
30=5d
30/5=d
6=d
Daughter's age = 6 years old
Suppose it takes x years to change the ratio from 5:1 to 3:1
In x years
Mother's age = 30+x
Daughter's age = 6+x
(30+x):(6+x)=3:1
(30+x)/(6+x)=3/1
(30+x)=3(6+x)
30+x=18+3x
30-18=3x-x
12=2x
12/2=x
6=x
x=6
In 6 years ratio between mother and daughter's will be 3:1
Check
30+6/6+6=3/1
36/12=3/1
3/1=3/1